OpenStudy (anonymous):

6. The navigator has to program the ship’s computer so that it will be prepared to intercept the satellite. To program it, the satelite’s path is set equal to the ship’s. The programmed equation is –4x – 2 = –x + 3. Explain to the navigator how the programmed equation can be graphed to find the point of intersection. Use complete sentences.

OpenStudy (anonymous):

The original problem is: Danielle is playing the newest app on her phone, Algebra City Defense. Earlier in the game, you and Danielle helped the Mathonauts defend the city against a giant, three-headed monster. After defeating the monster, the game advances to Level Two! You and Danielle must help the Mathonauts blast off to their Interstellar Headquarters. I chose y=5x+1 for my equation

OpenStudy (anonymous):

@wolfe8 ?

OpenStudy (wolfe8):

DANG I DID THIS BEFORE. Haha. Lemme find the link

OpenStudy (anonymous):

@wolfe8 lol okay

OpenStudy (wolfe8):

Ugh I did too many questions to find it. But basically you can separate the equation into two. One on the left side, one on the right side.

OpenStudy (anonymous):

so -4x-2 & -x+3?

OpenStudy (anonymous):

solve for x ?

OpenStudy (wolfe8):

Only if you want to find the interception point. But to sketch the paths, you will sketch y=-4x-2 and y=-4+3

OpenStudy (anonymous):

It says interception, so that's what I'm solving for I guess lol

OpenStudy (anonymous):

so the first one would be 6x? and the second would be -7?

OpenStudy (wolfe8):

Wait you're looking for the interception? Ah well then you just solve for x the equation before you separated them.

OpenStudy (anonymous):

I don't understand..

OpenStudy (wolfe8):

Well, if you're looking for the intersection point, you can solve for x in –4x – 2 = –x + 3 But since the question asked using graphs, that is why I told you to separate it into 2 straight line equations. Then you just graph/sketch the lines and see where they intersect.

OpenStudy (anonymous):

After I solve for x I'm lost. I'm sorry

OpenStudy (wolfe8):

Then you use the 2 equations we have to make 2 lines(paths)

OpenStudy (anonymous):

so after I solve for X and separate the two equations what do I do?

OpenStudy (wolfe8):

Draw the lines...

OpenStudy (anonymous):

its a flvs assignment so everything just has to be explained in words

OpenStudy (wolfe8):

Hehe well I can't give you the answer to that. Try this: http://www.mathsisfun.com/algebra/linear-equations.html

OpenStudy (anonymous):

thanks

OpenStudy (wolfe8):

You're welcome. Good luck.